# Lean classical-quantum hybrid neural network model for image classification

### v2

@@ Abstract
The integration of algorithms from quantum information with neural networks has enabled unprecedented advancements in various domains. Nonetheless, the application of quantum machine learning algorithms for image classification predominantly relies on traditional architectures such as variational quantum circuits. The performance of these models is closely tied to the scale of their parameters, with the substantial demand for parameters potentially leading to limitations in computational resources and a significant increase in computation time. In this paper, we introduce a Lean Classical-Quantum Hybrid Neural Network (LCQHNN), which achieves efficient classification performance with only four layers of variational circuits, thereby substantially reducing computational costs. We apply the LCQHNN to image classification tasks on public datasets and achieve a classification accuracy of 99.02% on the dataset, marking a 5.07% improvement over traditional deep learning methods. Under the same parameter conditions, this method shows a 75% and 70.59% improvement in training convergence speed on two datasets. Furthermore, through visualization studies, it is found that the model effectively captures key data features during training and establishes a clear association between these features and their corresponding categories. This study confirms that the employment of quantum algorithms enhances the model’s ability to handle complex classification problems.
@@ 1 Introduction
Neural Networks (NNs), as a vital machine learning tool, have achieved widespread success across various domains [1, 2, 3, 4, 5, 6]. Extensive research has demonstrated the formidable capabilities of neural networks in pattern recognition, feature extraction, and classification decision-making [7, 8, 9, 10]. However, the performance of traditional deep learning algorithms is intimately linked to the scale of the model and the size of the dataset. An increase in the scale of parameters can lead to constraints in computational resources and a significant escalation in computation time. Moreover, the optimization of hyperparameters presents a challenge, as there is currently no universally effective optimization method akin to gradient descent for hyperparameter optimization. When the parameter scale is large, the temporal cost of evaluating a set of hyperparameter configurations becomes exceedingly high. Recently, there has been a notable increase in interest regarding the integration of neural networks with quantum technologies [11, 12, 13, 14, 15, 16, 17]. Quantum algorithms possess a range of distinctive advantages that have the potential to significantly enhance both the performance and efficiency of traditional computational methods. By harnessing quantum parallelism, quantum-enhanced neural networks can navigate a broader search space within a reduced timeframe. Simultaneously, these networks can exploit quantum entanglement to establish correlations among complex data structures such as images, language representations, and time series data. Such correlations may facilitate more effective identification of underlying patterns within datasets by neural networks.
Variational quantum circuits (VQCs) are quantum algorithms designed to address optimization problems and machine learning tasks [18, 19, 20, 21]. The fundamental principle of VQCs is to identify the optimal solution by adjusting the parameters of the quantum circuit. The parameter count of VQCs is typically modest. This is because the design of VQCs aims to determine the optimal solution by adjusting a limited number of quantum circuit parameters, rather than relying on a large scale of parameters. By learning features undetectable by classical neural networks, VQCs can significantly enhance robustness against classical adversarial attacks, suggesting a potential quantum advantage in machine learning tasks [22, 23]. A quantum convolutional neural network (QCNN) has also been proposed for high-energy physics event classification [24], demonstrating faster convergence and higher test accuracy compared to CNNs [25]. Several studies have addressed issues such as gradient vanishing, robustness, and learnability [26, 27], thereby overcoming some of the inherent challenges. In recent years, several technology giants such as IBM, Google, Rigetti, Amazon Braket, and Microsoft Quantum Azure have deployed quantum computing hardware on cloud services. These machines are characterized as Noisy Intermediate-Scale Quantum (NISQ) devices. NISQ devices are typified by their possession of a moderate number of qubits and a susceptibility to errors due to the fragile nature of quantum states[28, 29, 30, 31, 32, 33, 34]. Despite these limitations, NISQ devices have demonstrated superior efficiency compared to classical computers in simulating complex quantum systems and performing certain types of computations[35, 36, 37, 38, 39].
On the other hand, although the participation of quantum algorithms can help models efficiently complete image classification tasks with limited computing resources, there are still many difficulties and limitations in specific applications, mainly from the following two aspects. The ‘curse of dimensionality’ is one of the main challenges in using quantum machine learning algorithms for image classification. Some image data have a huge spatial dimension, and as the data dimension increases, VQCs needs to process more qubits and perform more quantum gate operations, which can lead to exponential growth in computation time and resources. This computational limitation makes building a quantum image classification model that can meet practical application needs both time-consuming and expensive. On the other hand, complex quantum systems are highly sensitive to noise, and increasing the depth of VQCs can lead to more noise accumulation. The loss or error of quantum information can limit their performance in practical applications. Improving the efficiency of algorithms and ensuring the stability and reliability of quantum computing in image classification tasks remains a challenge.
@@
The development of effective quantum image classification models and the minimization of noise impact are critical research areas in machine learning. This paper introduces a lean classical-quantum hybrid neural network (LCQHNN) for image classification. We employ multi-channel convolution operators to extract nonlinear features from images, thereby reducing spatial dimensions while capturing essential data information. This enhances compatibility with current hardware capabilities. The streamlined four-layer variational quantum circuit (VQC) module is the present model can reduce quantum circuit complexity and mitigates noise sensitivity, thereby improving stability and reliability. Experimental results on public datasets demonstrate that LCQHNN significantly outperforms convolutional neural networks (CNNs) in terms of convergence speed and classification accuracy. Specifically, it achieves up to 75% and 70.59% faster training convergence on two datasets, while also achieving superior test set accuracy. In addition, we utilize Grad-CAM heat maps to visualize the prediction process of LCQHNN, revealing how it learns edge contours without prior knowledge. These visualizations enhance our understanding of the model’s decision-making process and validate its robustness in image recognition tasks.
@@ 2 Methodology
2.1 Quantum Machine Learning Algorithms
@@
Let’s first organize our theoretical framework. The quantum machine learning algorithm proposed in this study is a classical quantum classical process. We position the classical convolutional neural network at the front end of VQCs for feature extraction while reducing the spatial dimension to meet the quantum bit requirements of LCQHNN[40]. Although we have reduced the spatial dimension of the image, CNN maps the information that needs to be focused on to a higher feature dimension by increasing the number of channels, which is used to extract hidden features and improve accuracy[41]. Therefore, the features extracted by the CNN algorithm can be used as inputs for VQCs, where classical state data becomes quantum states. VQCs is the core of our algorithm, similar to the general approximation theorem in Artificial Neural Network. VQCs represents a circuit that can fit the objective function. During the training process, VQCs will repeat the feedforward process and compare the results with the true labels of the training data. By iteratively adjusting the parameters of the variational quantum circuit, the cost function is minimized in terms of the difference between the predicted label and the true label. Finally, the decoded measurement results provide the classification results of the input data, at which point the data state returns from quantum state to classical state.
@@
As shown in Figure 1, we construct the quantum algorithm model from five aspects to achieve image classification. Firstly, the feature extraction model of LCQHNN consists of two convolutional layers and two pooling layers. The convolutional layers are used to extract information to ensure reliable information transmission throughout the entire network architecture. Reducing the spatial dimension of the pooling layer can decrease the number of parameters and computational complexity in subsequent layers, making the network more efficient. In the second aspect, to ensure non-linear features, we added an activation layer composed of ReLU function between each convolutional layer and pooling layer. This work achieves nonlinear mapping of feature vectors. We assume that $f$ is the input image, $g$ is the convolution kernel, $(i,j)$ are the coordinates of the input image, and $(m,n)$ are the coordinates of the summation. So the convolution result should be $(f*g)(i,j)=\sum_{m=-\frac{k-1}{2}}^{\frac{k-1}{2}}\sum_{n=-\frac{k-1}{2}}^{\frac{k-1}{2}}f(m,n)g(i-m,j-n)$. Where k is the size of the convolution kernel, we set it to 3. The formula for ReLU is $max(0,x)$. We assume that $Window(i,j)$ represents a pooling window centered around $(i,j)$, and $P(i,j)$ is the value of the pooled output feature map at position $(i,j)$. So the extracted feature results are as follows[42].
@@
Thirdly, after feature extraction, we set up a transition model composed of fully connected networks, which reduces the dimensionality of features from 256 to 4 to accommodate quantum computing constraints. Due to the property of parameter sharing in convolutional neural networks, neurons in the network may develop complex co adaptation relationships. Therefore, we incorporate Dropout as a regularization technique into the module, which reduces co adaptation relationships among neurons by randomly discarding them. Intuitively described, it can be seen that each Dropout generates a different "sparse" network, which is equivalent to training many different networks and averaging their prediction results, which usually helps prevent overfitting of the training data by the network.
Fourthly, we use quantum encoding technology to transform the 4-dimensional feature vector into a 4-dimensional 4-bit quantum state, followed by a series of carefully designed single bit operations. Finally, by measuring the transition of quantum states to classical states, the classification results are obtained. We will provide a detailed description of the quantum algorithm in Section 2.2.
@@ 3.1.2 Datasets
In order to evaluate the performance of the proposed image classification model, the MNIST and Fashion MNIST datasets were selected for experiments.
MNIST (Modified National Institute of Standards and Technology database) is a large handwritten digit database created by Yann LeCun et al. in the 1990s[51]. It contains 60000 training samples and 10000 test samples, each of which is a 28x28 pixel handwritten digit image ranging from 0 to 9. The images in the MNIST dataset are grayscale and have been centered and standardized to ensure that the format of each image is consistent. Due to the large amount of computing resources and time required for training and validation processes, we have chosen a subset of MNIST, which includes 2048 balanced training samples and 512 balanced validation samples, respectively. We selected 1024 random samples as the test dataset to ensure the broad applicability of our method.
FashionMNIST is a dataset provided by Zalando (a German fashion e-commerce platform) as an alternative to MNIST[52], aiming to provide a similar dataset structure but containing clothing images rather than handwritten digits. Fashion MNIST contains 70000 images, divided into 10 categories (such as T-shirts, pants, shoes, etc.), with 7000 images per category and a size of 28x28 pixels. In this study, we trained on 2048 images labeled with two categories ("pants" and "shirt"), validated with 512 images, and tested with 1024 images.
@@ 3.1.3 Training Setting
3.2 Classification Analysis of MNIST and FashionMNIST
@@
We compared the LCQHNN against CNNs featuring fully connected modules with 4, 8, and 16 parameters, tracking the evolution of Accuracy throughout the training process. As can be seen from Figure 4(a), LCQHNN converges faster than a traditional cnn with 4 variable parameters, and the accuracy exceeds 99% by the sixth epoch, which is 75% quicker than its CNNs counterpart with the same parameters. Traditional CNNs necessitate a larger number of parameters to match this performance. Figure 4(b) displays the accuracy of the four models after 10 epochs of training. It is apparent that the model introduced in this paper performs on par with, and occasionally exceeds, the capabilities of sophisticated traditional neural network models on the digit classification task. This hints at the VQCs’s superior linear transformation capacity in comparison to the classical fully connected layers, allowing the model to concentrate on finer details, thereby attaining higher predictive precision. The data suggest that the LCQHNN possesses greater potential compared to conventional models.
@@
To further elucidate the differences in model performance, we trained four models on the more challenging FashionMNIST dataset and evaluated their classification accuracy on a separate test set, analogous to the structure of the previously discussed MNIST dataset. The experimental results are presented in Figure 5. Figure 5(a) illustrates the fluctuation in accuracy throughout the training process for all four models, with LCQHNN’s rapid convergence being particularly evident. The LCQHNN exceeded an accuracy threshold of 90% on the test set after only five epochs of training, representing a 70.59% improvement in convergence speed compared to traditional CNN models with an equivalent number of parameters. Even models with a larger array of parameters, such as CNN-8 and CNN-16, did not demonstrate such notable efficacy. This phenomenon can be attributed to VQCs’s adeptness at exploring high-dimensional feature spaces, facilitating quicker comprehension and adaptation to the underlying characteristics of the data. Figure 5(b) delineates each models precision on the FashionMNIST test set after ten epochs these outcomes substantiate that within an identical constrained training period, LCQHNN achieves superior classification precision, corroborating that incorporating quantum attributes significantly enhances classification proficiency within a reduced training duration.
@@
Table 1: Comparison of classification accuracy on public datasets.The table summarizes the classification accuracy of four models trained on the MNIST and FashionMNIST datasets for 10, 25, and 50 epochs, highlighting the unique highest accuracy within each group. If two or more models have the same accuracy, no emphasis is given.

### v3

@@ Abstract
The integration of algorithms from quantum information with neural networks has enabled unprecedented advancements in various domains. Nonetheless, the application of quantum machine learning algorithms for image classification predominantly relies on traditional architectures such as variational quantum circuits. The performance of these models is closely tied to the scale of their parameters, with the substantial demand for parameters potentially leading to limitations in computational resources and a significant increase in computation time. In this paper, we introduce a Lean Classical-Quantum Hybrid Neural Network (LCQHNN), which achieves efficient classification performance with only four layers of variational circuits, thereby substantially reducing computational costs. Our experiments demonstrate that LCQHNN achieves 100%, 99.02%, and 85.55% classification accuracy on MNIST, FashionMNIST, and CIFAR-10 datasets. Under the same parameter conditions, the convergence speed of this method is also faster than that of traditional models. Furthermore, through visualization studies, it is found that the model effectively captures key data features during training and establishes a clear association between these features and their corresponding categories. This study confirms that the employment of quantum algorithms enhances the model’s ability to handle complex classification problems.
@@ 1 Introduction
Neural Networks (NNs), as a vital machine learning tool, have achieved widespread success across various domains [1, 2, 3, 4, 5, 6]. Extensive research has demonstrated the formidable capabilities of neural networks in pattern recognition, feature extraction, and classification decision-making [7, 8, 9, 10]. However, the performance of traditional deep learning algorithms is intimately linked to the scale of the model and the size of the dataset. An increase in the scale of parameters can lead to constraints in computational resources and a significant escalation in computation time. Moreover, the optimization of hyperparameters presents a challenge, as there is currently no universally effective optimization method akin to gradient descent for hyperparameter optimization. When the parameter scale is large, the temporal cost of evaluating a set of hyperparameter configurations becomes exceedingly high. Recently, there has been a notable increase in interest regarding the integration of neural networks with quantum technologies [11, 12, 13, 14, 15, 16, 17]. Quantum algorithms possess a range of distinctive advantages that have the potential to significantly enhance both the performance and efficiency of traditional computational methods[18, 19, 20]. For instance, Liu et al. proposed a hybrid quantum-classical convolutional neural network (QCCNN), which employs parametric quantum circuits to replace classical convolutional kernels, enabling the model to better extract patterns from complex data distributions[21]. Similarly, A. Senokosov et al. introduced parallel quantum layers (HQNN-Parallel) and quantum convolutional layers (HQNN-Quanv), demonstrating that hybrid models can outperform classical counterparts even with fewer trainable parameters[22]. However, the same study also highlights that having fewer trainable parameters does not inherently guarantee more efficient execution in quantum machine learning (QML) models compared to classical ones. This limitation arises from the significantly higher computational overhead associated with quantum model training and inference.
Variational Quantum Circuits (VQCs) constitute a class of quantum algorithms designed to solve optimization and machine learning problems through tunable parameterized quantum circuits [23, 24, 25, 26]. Their core mechanism relies on optimizing a limited set of quantum gate parameters (typically maintaining controllable parameter scales) rather than relying on large-scale parameter accumulation to approximate optimal solutions. VQCs extract feature patterns through quantum entanglement hierarchies that are undetectable by classical neural networks [27], thereby significantly enhancing robustness against classical adversarial attacks and providing theoretical foundations for practical quantum advantages in machine learning [28]. Recent advancements in quantum machine learning have witnessed breakthroughs in two key technological pathways. On one hand, in the realm of hybrid quantum-classical feature engineering, Chen et al. [29] innovatively combined the multi-scale entanglement renormalization Ansatz with box-counting fractal feature extraction to construct a Quantum Convolutional Neural Network (QCNN). This network has demonstrated superior accuracy compared to classical Convolutional Neural Networks (CNNs) in medical imaging classification tasks for breast cancer. Moreover, the universality of this quantum feature engineering approach has been further validated in the field of high-energy physics. In particle collision event classification tasks [30], Chen et al. found that, with a comparable number of parameters, QCNNs outperformed classical CNNs [31] in both test accuracy and learning speed. On the other hand, to address the resource constraints of Noisy Intermediate-Scale Quantum (NISQ) devices, Zhou et al. proposed a hybrid quantum-classical Generative Adversarial Network (QGAN) [32]. This network innovatively employs image remapping techniques to transform complex multi-modal distributions into single-modal representations that are processable by quantum systems. The scheme successfully achieved high-quality reconstruction of MNIST and Fashion-MNIST images, marking a significant step toward the real-world deployment of quantum technologies.
Although the participation of quantum algorithms can help models efficiently complete image classification tasks with limited computing resources, there are still many difficulties and limitations in specific applications [33, 34], mainly from the following two aspects. The ’curse of dimensionality’ poses a significant challenge in implementing quantum machine learning algorithms for image classification tasks [35, 36, 37]. Image data typically exhibits extremely high spatial dimensions, and encoding such data into quantum states requires substantial quantum bit resources. Specifically, a system composed of n qubits has a state space dimension of $2^{n}$, which implies that the quantum resources required for VQCs grow exponentially with increasing data dimensionality, far exceeding the processing capabilities of current NISQ hardware. Meanwhile, complex quantum systems are highly sensitive to noise, and increasing the depth of VQCs can lead to more noise accumulation. For instance, Wang et al. (2021)[38] demonstrated that noise in NISQ devices induces barren plateaus in the training landscape (i.e., vanishing gradients). As the circuit depth of VQAs scales linearly with the number of qubits $n$, the parameter gradients vanish exponentially with $n$. In a parallel study, Aidan Pellow-Jarman et al. investigated the optimal circuit depth for the quantum approximate optimization algorithm (QAOA) and revealed that each additional layer introduces hardware noise [39]. Their experiments showed that task performance peaks at approximately six layers. Beyond this critical depth, noise accumulation dominates, leading to a gradual performance degradation. This noise-depth trade-off can be extended to high-dimensional tasks such as image classification. These studies indicate that the noise characteristics of current NISQ devices severely limit the scalability of quantum algorithms. Therefore, how to enhance algorithmic efficiency while ensuring the stability and reliability of quantum computing in image classification remains a critical challenge to be addressed.
@@
The development of effective quantum image classification models and the minimization of noise impact are critical research areas in machine learning. To realize quantum enhancement of image classification with limited hardware resources, this paper introduces a lean classical-quantum hybrid neural network (LCQHNN) for image classification. We employ classical multi-channel convolutional operators for nonlinear feature extraction from images. Compared to quantum convolution kernels, this approach offers superior hardware compatibility, ensuring better alignment with contemporary computing architectures. Meanwhile, we introduce a streamlined four-layer variational quantum circuit (VQC) module that effectively reduces quantum circuit complexity and mitigates noise sensitivity. By minimizing circuit depth and optimizing parameter sharing, our approach enhances stability and reliability, avoiding the decoherence errors and resource overhead often associated with parallel quantum circuits. Experimental results on public datasets demonstrate that LCQHNN significantly outperforms convolutional neural networks (CNNs) in terms of convergence speed and classification accuracy. Specifically, it achieved training convergence speeds that were 75% and 70.59% faster than those of CNNs with the same parameters on the MNIST and FashionMNIST datasets, while also achieving excellent accuracy on the test sets. In addition, we utilize Grad-CAM heat maps to visualize the prediction process of LCQHNN, revealing how it learns edge contours without prior knowledge. These visualizations enhance our understanding of the model’s decision-making process and validate its robustness in image recognition tasks.
@@ 2 Methodology
2.1 Introduction to the LCQHNN Algorithm
@@
Let’s first organize our theoretical framework. The quantum machine learning algorithm proposed in this study is a classical quantum classical process. We position the classical convolutional neural network at the front end of VQCs for feature extraction while reducing the spatial dimension to meet the quantum bit requirements of LCQHNN[40]. Although we have reduced the spatial dimension of the image, CNN maps the information that needs to be focused on to a higher feature dimension by increasing the number of channels, which is used to extract hidden features and improve accuracy[41]. Therefore, the features extracted by the CNN algorithm can be used as inputs for VQCs, where classical state data becomes quantum states. VQCs is the core of our algorithm, similar to the general approximation theorem in Artificial Neural Network. VQCs represents a circuit that can fit the objective function. During the training process, VQCs will repeat the feedforward process and compare the results with the true labels of the training data. By iteratively adjusting the parameters of the variational quantum circuit, the cost function is minimized in terms of the difference between the predicted label and the true label. Finally, the measurement results of the four qubits yield a set of 4-bit classical states. These classical data are then processed through a fully connected neural network for linear transformation, and subsequently normalized via the softmax activation function to produce a probabilistic distribution over all target classes.
Prior to feature extraction, the input images undergo standardized preprocessing to ensure compatibility with the LCQHNN architecture. For the dataset used in the experiment mentioned in this paper, raw images are uniformly processed as follows.
  • Grayscale Normalization:
    • For grayscale datasets (MNIST/FashionMNIST), pixel values are linearly scaled from $[0,255]$ to $[0.0,1.0]$:
      $I_{\text{norm}}=I_{\text{raw}}/255.0$
    • For RGB datasets (CIFAR-10), each channel is independently normalized:
      $I^{(c)}_{\text{norm}}=I^{(c)}_{\text{raw}}/255.0,\quad c\in\{R,G,B\}$
  • Spatial Standardization:
    • MNIST/FashionMNIST images are resized to $28\times 28$ pixels
    • CIFAR-10 images are resized to $32\times 32$ pixels
  • Tensor Conversion:
    • MNIST/FashionMNIST: $[1,28,28]$ (single channel)
    • CIFAR-10: $[3,32,32]$ (RGB channels)
@@
As shown in Figure 1, we construct the quantum algorithm model from five aspects to achieve image classification. Firstly, the feature extraction model of LCQHNN consists of two convolutional layers and two pooling layers. The convolutional layers are used to extract information to ensure reliable information transmission throughout the entire network architecture. Reducing the spatial dimension of the pooling layer can decrease the number of parameters and computational complexity in subsequent layers, making the network more efficient. In the second aspect, to ensure non-linear features, we added an activation layer composed of ReLU function between each convolutional layer and pooling layer. This work achieves nonlinear mapping of feature vectors. We assume that $f$ is the input image, $g$ is the convolution kernel, $(i,j)$ are the coordinates of the input image, and $(m,n)$ are the coordinates of the summation. So the convolution result should be $(f*g)(i,j)=\sum_{m=-\frac{k-1}{2}}^{\frac{k-1}{2}}\sum_{n=-\frac{k-1}{2}}^{\frac{k-1}{2}}f(m,n)g(i-m,j-n)$. Where k is the size of the convolution kernel, we set it to 3. The formula for ReLU is $max(0,x)$. We assume that $Window(i,j)$ represents a pooling window centered around $(i,j)$, and $P(i,j)$ is the value of the pooled output feature map at position $(i,j)$. So the extracted feature results [42] are as follows.
@@
Thirdly, after feature extraction, we set up a transition model composed of fully connected networks, which reduces the dimensionality of features from 256/400 to 4 to accommodate quantum computing constraints, the fully connected input dimension depends on the size of the initial input image. Due to the property of parameter sharing in convolutional neural networks, neurons in the network may develop complex co adaptation relationships. Therefore, we incorporate Dropout as a regularization technique into the module, which reduces co adaptation relationships among neurons by randomly discarding them. Intuitively described, it can be seen that each Dropout generates a different "sparse" network, which is equivalent to training many different networks and averaging their prediction results, which usually helps prevent overfitting of the training data by the network.
Fourthly, we use quantum encoding technology to transform the 4-dimensional feature vector into a 4-dimensional 4-bit quantum state, followed by a series of carefully designed single bit operations. Finally, the quantum measurement results are processed through a fully connected network and softmax function to obtain the classification outcome. We will provide a detailed description of the quantum algorithm in Section 2.2.
@@ 3.1.2 Datasets
In order to evaluate the performance of the proposed image classification model, the MNIST, Fashion MNIST, and CIFAR-10 datasets were selected for experiments.
MNIST (Modified National Institute of Standards and Technology database) is a large handwritten digit database created by Yann LeCun et al. in the 1990s [51]. It contains 60,000 training samples and 10,000 test samples, each of which is a $28\times 28$ pixel handwritten digit image ranging from 0 to 9. The images in the MNIST dataset are grayscale and have been centered and standardized to ensure that the format of each image is consistent. Due to the large amount of computing resources and time required for training and validation processes, we have chosen a subset of MNIST, which includes 2048 balanced training samples and 512 balanced validation samples, respectively ("0" and "1"). We selected 1024 random samples as the test dataset to ensure the broad applicability of our method.
FashionMNIST is a dataset provided by Zalando (a German fashion e-commerce platform) as an alternative to MNIST [52], aiming to provide a similar dataset structure but containing clothing images rather than handwritten digits. FashionMNIST contains 70,000 images, divided into 10 categories (such as T-shirts, pants, shoes, etc.), with 7000 images per category and a size of $28\times 28$ pixels. In this study, we trained on 2048 images labeled with two categories ("pants" and "shirt"), validated with 512 images, and tested with 1024 images.
Additionally, we used the CIFAR-10 dataset [53], which consists of 60,000 $32\times 32$ color images divided into 10 classes, with 6,000 images per class. For our binary classification task, we selected two classes ("airplane" and "car") from CIFAR-10. The training set included 2048 images, the validation set included 512 images, and the test set included 1024 images. This selection allowed us to further evaluate the model’s performance on more complex, color images while maintaining a balanced dataset.
@@ 3.1.3 Training Setting
3.2 Classification Analysis of Multiple Datasets
@@
Table 1: Model Architecture Comparison
Model Architecture Description Number of parameters
LCQHNN
Variational Quantum Circuit (VQC)
$\rightarrow$ FC(4$\rightarrow$2)
$4+(4\times 2)\\ =8_{\text{weights}}\end{aligned}$
CNN-4
FC(4$\rightarrow$1)
$\rightarrow$ FC(1$\rightarrow$2)
$(4\times 1)+(1\times 2)\\ =6_{\text{weights}}\end{aligned}$
CNN-8
FC(4$\rightarrow$2)
$\rightarrow$ FC(2$\rightarrow$2)
$(4\times 2)+(2\times 2)\\ =12_{\text{weights}}\end{aligned}$
CNN-16
FC(4$\rightarrow$4)
$\rightarrow$ FC(4$\rightarrow$2)
$(4\times 4)+(4\times 2)\\ =24_{\text{weights}}\end{aligned}$
We compared the LCQHNN against CNNs featuring fully connected modules with 4, 8, and 16 parameters, tracking the evolution of Accuracy throughout the training process. The classification module differences between models are shown in the Table 1, where the numerical suffix (4/8/16) denotes the parameter count of the first FC layer in the classical models. As can be seen from Figure 4(a), LCQHNN converges faster than a traditional cnn with 4 variable parameters, and the accuracy exceeds 99% by the sixth epoch, which is 75% quicker than its CNNs counterpart with the same parameters. Traditional CNNs necessitate a larger number of parameters to match this performance. Figure 4(b) displays the accuracy of the four models after 10 epochs of training. It is apparent that the model introduced in this paper performs on par with, and occasionally exceeds, the capabilities of sophisticated traditional neural network models on the digit classification task. This hints at the VQCs’s superior linear transformation capacity in comparison to the classical fully connected layers, allowing the model to concentrate on finer details, thereby attaining higher predictive precision. The data suggest that the LCQHNN possesses greater potential compared to conventional models.
@@
Based on the previous experimental results, we trained four models on the more challenging FashionMNIST dataset and evaluated their classification accuracy on a separate test set, analogous to the structure of the previously discussed MNIST dataset. The experimental results are presented in Figure 5. Figure 5(a) illustrates the fluctuation in accuracy throughout the training process for all four models, with LCQHNN’s rapid convergence being particularly evident. The LCQHNN exceeded an accuracy threshold of 90% on the test set after only five epochs of training, representing a 70.59% improvement in convergence speed compared to traditional CNN models with an equivalent number of parameters. Even models with a larger array of parameters, such as CNN-8 and CNN-16, did not demonstrate such notable efficacy. This phenomenon can be attributed to VQCs’s adeptness at exploring high-dimensional feature spaces, facilitating quicker comprehension and adaptation to the underlying characteristics of the data. Figure 5(b) delineates each models precision on the FashionMNIST test set after ten epochs these outcomes substantiate that within an identical constrained training period, LCQHNN achieves superior classification precision, corroborating that incorporating quantum attributes significantly enhances classification proficiency within a reduced training duration.
@@
Table 2: Comparison of classification accuracy on public datasets.
@@
To further validate the generalizability of the LCQHNN model, we systematically compared it with current mainstream image classification architectures in benchmark tests. Among them, HQNN (Hybrid Quantum Neural Network)[50], as a similar hybrid architecture, is collaboratively constructed with classical convolutional modules and variational quantum circuits (VQCs), and its design paradigm originates from the cutting-edge research of the Qiskit machine learning team [^1] Code available at: https://qiskit-community.github.io/qiskit-machine-learning/tutorials/05_torch_connector.html. QNN (Quantum Neural Network)[55], on the other hand, is a fully quantum circuit-based end-to-end architecture that does not integrate any classical computational units [^2] Code available at: https://quict-docs.readthedocs.io/aa/latest/tutorials/algorithm/QNN. By introducing the CIFAR-10 datasets, which contains complex textural features, this experiment aims to verify whether LCQHNN can maintain its architectural advantages when faced with visual tasks exhibiting significant cross-domain feature variations. As shown in Table 3, the experimental results demonstrate that in the binary classification task on CIFAR-10, our proposed LCQHNN achieves the best performance (85.55%) compared to CNN-16, HQNN, and QNN. Except for QNN, all other models achieve perfect accuracy on MNIST, highlighting the necessity of classical-quantum hybridization. On FashionMNIST, although HQNN achieves slightly higher accuracy than LCQHNN (99.12% vs. 99.02%), this result does not overshadow LCQHNN’s innovative architecture and its exceptional performance on CIFAR-10. The minor difference on FashionMNIST may be attributed to the dataset’s lower complexity, which aligns well with HQNN’s design. In contrast, LCQHNN’s ability to excel in both simple and highly complex tasks underscores its potential as a versatile and future-proof solution for advanced image classification challenges.
Table 3: Performance comparison of LCQHNN with classical, quantum, and hybrid models on benchmark datasets.
Model MNIST FashionMNIST CIFAR-10
LCQHNN (our) 100% 99.02% 85.55%
CNN-16 100% 98.24% 83.89%
HQNN 100% 99.12% 83.79%
QNN 85.93% 78.13% 60.93%
Figure 6 illustrates the confusion matrices of LCQHNN across three datasets, highlighting its classification performance on varying image complexities. In the MNIST binary task, the model achieved 100% accuracy with zero misclassifications: all 512 negative samples ("0") and 512 positive samples ("1") were perfectly categorized. For Fashion-MNIST, the model attained 99.02% accuracy, with minor errors including 7 false positives (pants misclassified as shirts) and 3 false negatives (shirts misclassified as pants). On the more challenging CIFAR-10 subset, LCQHNN achieved 85.55% accuracy in distinguishing airplanes from cars. While correctly identifying 443 airplanes (TN) and 433 cars (TP), it exhibited 69 false positives (airplanes misclassified as cars) and 79 false negatives (cars misclassified as airplanes), reflecting the dataset’s inherent complexity in fine-grained feature discrimination. These results demonstrate LCQHNN’s robustness in simple classification tasks and its competitive capability in complex scenarios, with errors primarily arising from subtle inter-class variations in advanced vision tasks.
Figure 6: The confusion matrices of LCQHNN on two distinct datasets. In the binary classification task of MNIST (left), the model achieved 100% accuracy. On the Fashion-MNIST dataset (middle) and CIFAR-10 dataset, minor classification errors were observed.
Although LCQHNN demonstrates promising performance on the Fashion-MNIST and CIFAR-10 datasets, its practical application is subject to several critical limitations. First, the experiments were conducted only on small-scale datasets ($\leq$70k samples), and the model’s performance on larger-scale tasks (e.g., ImageNet) remains untested, potentially requiring architectural adaptations to handle increased data complexity. Second, the current evaluations rely on classical quantum simulators, which cannot fully replicate the noise characteristics of real quantum hardware. This discrepancy may lead to deviations between simulated results and actual performance in real-world quantum environments. Finally, it should be pointed out that our current model design is only for binary classification tasks, which to some extent limits its applicability to complex multi classification scenarios. We will investigate this issue in future work.